OpenStudy (anonymous):

Find the unit vector having the same direction as v. v = -5i + 12j

OpenStudy (anonymous):

so far I have:

OpenStudy (anonymous):

\[v = -5i+12j\]\[||v|| = \sqrt{(-5)^2 + 12^2}\]\[||v|| = \sqrt{25 + 144}\]\[||v|| = \sqrt{169}\]\[||v|| = 13\] So: \[u = \frac{ -5i+12j }{ 13 }\]\[u = \frac{ -5 }{ 13 }+\frac{ 12 }{ 13}\] Therefore: \[||u|| = \sqrt{(\frac{ -5 }{ 13})^2+(\frac{ 12}{ 13})^2}\]\[||u|| = \sqrt{\frac{ 25 }{ 169 }+\frac{ 144 }{ 169 }}\]\[||u|| = \sqrt{\frac{ 169 }{ 169}}\]\[||u|| = 1\]

OpenStudy (amistre64):

we want to scale it by some number: k, such that (5k)^2 + (12k)^2 = 1

OpenStudy (amistre64):

you did fine

OpenStudy (anonymous):

I know the answer is not, \[||u|| = 1\] is it \[u = \frac{ -5 }{ 13 }i+\frac{ 12 }{ 13 }j\]?

OpenStudy (anonymous):

i can do the math, but I am not 100% of which answer it wants

OpenStudy (amistre64):

its the scaled version of u, the divide by 13 stuff

OpenStudy (anonymous):

so, finding the magnitude of u is just extra?

OpenStudy (amistre64):

its not asking for the length of the unit vector its asking for the vector itself, that has a length of 1 and is in the same direction as u well, any positive scalar of u is in the same directions as u

OpenStudy (amistre64):

no, u is 13 units long ... we want it 1 unit long 13/13 = 1

OpenStudy (anonymous):

awesome, once again thank you! I doubt myself a lot.

OpenStudy (amistre64):

consider a triangle 

OpenStudy (amistre64):

similar triangles 1:13 a:5 b:12 what are a and b?

OpenStudy (anonymous):

they should be a ratio of the sides they parallel, no?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1=a^2 + b^2

OpenStudy (amistre64):

we scaled 13 down to 1 ... how? divide by 13 a = 5/13 b = 12/13 since 1=13/13

OpenStudy (anonymous):

ignore what i said just now. i had a stroke, not a stroke of genius xD

OpenStudy (amistre64):

a:b:c 5:12:13 c = 1 tho, divide by 13 to scale the ratio 5/13 : 12/13: 13/13

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

Thank you for your over the top help!

OpenStudy (amistre64):

yep in short, a unit vector is just the vector scaled so its length is 1

OpenStudy (anonymous):

yay for learning!